what is 1/25 as a percent and how did you get it? Thanks
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

what is 1/25 as a percent and how did you get it?
---
1/25 = 4/100 = 0.04
----
0.04*0.01
= 0.0004
= 0.04/100
= 0.04%
============
Cheers,
Stan H.
==============================================

Answer by Alan3354(69443)

1/25 = 4/100
= 4%